OpenStudy (anonymous):

A mass of 10.0 kilograms is completely converted into energy. What is this amount of energy in joules? (1 J = 1 ) 3.33 × 108 J 1.11 × 10-8 J 3.00 × 1016 J 9.00 × 1017 J

OpenStudy (anonymous):

(1 J = 1 kg * m^2 / s^2)

OpenStudy (anonymous):

E = m * c^2 = 10.00 Kg * (299,792,458 m/s)^2 = 2997,92 Joules

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

The equation that explain you the realtion between the mass and energy is given by E = m * c^2. Where E is energy m: is the mass of the object c: is the speed of light, which is a costant (299,792,458 m/s)

OpenStudy (photon336):

\[1 joule = 1 newton \] Newton \[\frac{ m ^{2} }{ s ^{2} }*kg\] those are your units.

OpenStudy (photon336):

newton meter***